Ticket Number: HEA-NOC/20040616-3 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: LanComms
Ticket Opened: 20040616 10:18 UTC+1 Problem Start: 20040616 07:10 UTC+1
Ticket Closed: 20040723 14:14 UTC+1 Problem End: 20040616 07:12 UTC+1
Site/line: DCU, DIAS, DIT, MLE, NCAD, NCIRL, NUIM, RCSI, TCD, UCD
Router logs report VIP crash on a router in Brooklawn House responsible for
ATM backup links for clients.
Backup connections to affected clients unavailable for the duration of the
20040616 10:14 AH Report to Lancomms.
20040617 09:48 AH A software bug is suspected as the cause of this
problem. Lancomms are investigating a suitable upgrade path if applicable.
An IOS upgrade is provisionally scheduled for the maintenance window at
08:00 on Wednesday 23/6/04. This upgrade would affect backup links only.
20040622 16:46 BN Due to further complications with this router the IOS
upgrade will not be performed on Wednesday 23/06/04. A schedule of work
will be decided upon and this ticket will be updated accordingly.
20040622 17:12 BN Some configuration changes, as suggested by Cisco TAC,
are now planned for Thursday morning. This should be non-service
affecting. These changes would ease some current problems in preparation
for further work.
20040624 09:33 BN Software config changes have been made. Effects will be
20040628 16:59 WD issues arise with teleute. WD calls LANCOMMS awaiting
20040628 17:07 WD: emails John & [log in to unmask] with syslog
20040707 16:00 BN Further work, as suggested by Cisco TAC, will be carried
out between 08:00 and 10:00 on Thursday 15 July. Router will be powered
off during work, but no direct impact on customers is expected as this is
a backup only. Work plan is as follows:
1) Power down router.
2) Remove cards from slots 1 & 5.
3) Remove VIP2-50 from slot0
5) Strip cards from VIP2-50
6) Place cards into replacement VIP2-50 or VIP4-80 if possible.
7) Insert replacement VIP card.
8) Power up router and test.
20040713 09:30 JC tested VIP4-80 in testlab, was recognised by router,
recognised modules when installed in VIP.
20040713 10:34 RD tcd-atmbak (8:21-8.23) and diasbak (3:39-3:42) lines
were down on Teleute this morning. Both lines recovered within minutes.
20040715 09:08 JC and CW removed VIP2-50s and installed VIP4-80 along with
the ATM and FastEthernet modules in Teleute. OSPF came up. Monitoring for
20040723 14:13 JC Teleute appears to be working fine. Closing ticket.
This ticket can be monitored at https://www.hea.net/mousetrap/view/20040616-3